OpenStudy (anonymous):

The number of traffic violations in a city increased from 12 to 16 in a month. What is the percent of increase in the number of traffic violations for that month? Round to the nearest percent, if necessary. A. 20% B. 25% C. 33% D. 75%

563blackghost (563blackghost):

Increase is from 12 to 16 so an increase of 4. As a percentege of 12 the calculation is:- 4/12 x 100 = 400/12 = 33.33 recurring so to nearest percent would be 33% if this was helpful please leave a medal xD